OpenStudy (vladthepro):

Identify the 25th term of the arithmetic sequence 2, 1 and 3 over 5 , 1 and 1 over 5 ... −7 3 over 5 −8 −12 2 over 5 −13

OpenStudy (mathmale):

Please, Vlad, write that in mathematical symbols. You might get faster results that way.

OpenStudy (fortytherapper):

Right, I have no idea if it's this: \[-7 \frac{ 3 }{ 5 }, -8,-12\frac{ 2 }{ 5 },-13\]

OpenStudy (vladthepro):

symbols for solutions?

OpenStudy (mathmale):

arithmetic sequence 2, 1 and 3 over 5 , 1 and 1 over 5 ... |

OpenStudy (vladthepro):

sorry. I'm lagging

OpenStudy (vladthepro):

Q: http://prntscr.com/aikwjq @mathmale

OpenStudy (mathmale):

thank you for posting that illustration. SO much clearer! the formula for the nth term of an arith sequence is L=A+(n-1)D, ...where A is the first ter, n is the nyumber of terms, D is the difference (gap) between terms, and L is the last term.

OpenStudy (vladthepro):

Ohh! I see. Makes sense. Thanks buddy! :)